OpenStudy (mirandaregina):

How do I reduce 8+4√3/8 to its lowest term? According to my calculator, the answer is 2+√3/2, but I would like to know how to get this answer without depending on my calculator.

OpenStudy (mirandaregina):

it's 8+4 square root of 3 over 2

OpenStudy (anonymous):

is the square root over 3/2 or just the 3?

OpenStudy (mirandaregina):

@malapaka95 just the 3. 8+4 square root of 3/2

OpenStudy (anonymous):

according to your explanation its \[8 + \sqrt{3}/2\]

OpenStudy (mirandaregina):

@malapaka95 I'm trying to reduce  sorry if I confused you :)

OpenStudy (anonymous):

ohhhh haha

OpenStudy (anonymous):

well this is just : \[8/8 +( 4/8 \times \sqrt{3}) = 1 +( 1/2 * \sqrt{3}) = 1 + \sqrt{3}/2\]

OpenStudy (mirandaregina):

@malapaka95 ok, thank you so much!